OpenStudy (anonymous):

2x²+15 x+18 solve by completing the square

OpenStudy (anonymous):

is this deriving?if it is then the answer is 4x+15

OpenStudy (anonymous):

start with \[2x^2+15x=-18\] divide by 2 and get \[x^2+\frac{15}{2}x=-9\] then complete the square by taking half of \(\frac{15}{2}\) which is \(\frac{15}{4}\) square and add to both sides to get \[(x+\frac{15}{4})^2=-9+(\frac{15}{4})^2\] now some arithmetic

OpenStudy (anonymous):

\[-9+(\frac{15}{4})^2=\frac{81}{16}\] so you get \[(x+\frac{15}{4})^2=\frac{81}{16}\] take the root and get \[x+\frac{15}{4}=\frac{9}{4}\] or \[x+\frac{15}{4}=-\frac{9}{4}\] then subtract \(\frac{15}{4}\) in both equations to solve for \(x\)

OpenStudy (anonymous):

shouldnt their be two answwers?

OpenStudy (anonymous):

i know one of them is -6, and rge other should be -3/2..... but how do i get that??

OpenStudy (anonymous):

Maybe u want IT=(2x+3)*(x+6)

OpenStudy (anonymous):

huh?? if your talking about the equation.. i cant change that! ;P

OpenStudy (anonymous):

yes there are two answers \[x+\frac{15}{4}=\frac{9}{4}\] \[x=\frac{9}{4}-\frac{15}{4}=-\frac{6}{4}=-\frac{3}{2}\] is one of them

OpenStudy (anonymous):

\[x+\frac{15}{4}=-\frac{9}{4}\] \[x=-\frac{9}{4}-\frac{15}{4}=-\frac{24}{4}=-6\]is the other one

OpenStudy (anonymous):

thank you sooooooooo much your a life saver!!;)